In lab, you diluted your unknown protein solution by combining 14 LaTeX: \mu\text{L}?L of unknown protein solution with enough PBS buffer to obtain a total volume of 100 LaTeX: \mu\text{L}?L, then added 1000 LaTeX: \mu\text{L}?L Bradford reagent. After waiting ten minutes, the absorbance at 595 nm was 0.206. Using the standard curve from the previous question, calculate the total concentration of protein (in LaTeX: \frac{\mu\text{g}}{\text{mL}\:}?gmL) in the original sample.

Slope = 0.02371    Intercept = 0.01132    r2 = 0.973
Answer #1

Assuming that absorption is on y-axis and concentration on x-axis,

A = slope * Concentration + Intercept

0.206 = 0.02371 C + 0.01132

0.206 - 0.01132 = 0.02371 C

0.19468 = 0.2371 C

C = 0.19468/0.2371 = 0.8211

To calculate the original concentration in 14uL of solution, we need not worry about 1000uL of bradford reagent because that was added to each of the standard solutions. However, we have diluted our sample from 14uL to 100uL. So, the original concentration will be

original C = 0.8211 * 100/14 = 5.865 in the units of the concentration of standard solutions.